OpenStudy (anonymous):

Tammy is at the dentist's office waiting on her appointment. She notices that the 6-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 1:20 to 1:55? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches? Part 4: What is the coordinate point associated with this radian measure?

OpenStudy (anonymous):

@satellite73 you have some time to help me answer these four parts? :/

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well 60 minutes = 2 pi radians on the clockface so 1 minute = 2pi/60 radians then 1:20 to 1:55 = 35 mins 35 minutes = 35 * 2pi/60

OpenStudy (campbell_st):

(b) the arc length is \[l = r \theta\] you have theta from the previous part and the minute hand is 6 inches... so r = 6

OpenStudy (campbell_st):

(C) you will have to use arc length again \[5\pi = 6 \times \theta\] find theta this will give the angle in radian hope it helps

OpenStudy (anonymous):

by c you mean 3?

OpenStudy (anonymous):

how do i find theta?

OpenStudy (campbell_st):

yep

OpenStudy (campbell_st):

and divide both sides of the equation by 6

OpenStudy (anonymous):

theta=5pi/6

OpenStudy (campbell_st):

yes

OpenStudy (anonymous):

for #2 how do i find what there asking for? l=6*theta

OpenStudy (anonymous):

how to find the lenght?

OpenStudy (campbell_st):

well 1 minute = 2pi/60 so the 35 minutes through with the minute hand travels form and angle 35*2pi/60 or 7pi/6 thats the value of theta

OpenStudy (anonymous):

so.l=6*7pi/6?

OpenStudy (campbell_st):

yes

OpenStudy (anonymous):

can that be simplified?

OpenStudy (campbell_st):

well there is an obvious common factor to cancel

OpenStudy (anonymous):

l=7pi?

OpenStudy (campbell_st):

yes

OpenStudy (anonymous):

thank you for being patient... would that answer #2

OpenStudy (campbell_st):

its is

OpenStudy (anonymous):

ok going back to question one..35*2pi/60=219.8?

OpenStudy (campbell_st):

no... leave the answer in terms of pi.... you are working in radian its \[\frac{35 \times 2\pi}{60}\] simplify it

OpenStudy (anonymous):

can we simplify 7*2pi/12 ?

OpenStudy (campbell_st):

sure its 14pi/12 they are both even...

OpenStudy (anonymous):

7pi/6 then?

OpenStudy (anonymous):

now we have answered #1 and 2

OpenStudy (campbell_st):

yes

OpenStudy (campbell_st):

if you go back and read the posts you have the answers to 1, 2, and 3

OpenStudy (anonymous):

yes i meant 3 also :) so how will we find #4

OpenStudy (campbell_st):

 just pick up a calculator adnd find the values of x and y